A wire of length 0.50 m is fixed horizontally between two supports separated by 0.50 m. When a

mass of 8 kg hangs from the middle of the wire, the mid-point sags by 1 cm. The diameter of the

wire is 2.8 mm. Calculate the Young’s modulus of the wire.


1
Expert's answer
2021-10-31T18:13:26-0400

Find the change in length. First, find the stretched length:


"L_s=2\sqrt{(0.5/2)^2+0.01^2}=0.5004\text{ m}."

Find the tension in the wire:


"2T\sin\arctan\frac{L_\text{sag}}{L/2}=mg,\\\space\\
T=981\text{ N}."

Strain:


"\epsilon=\frac{L_s-L}{L}=\frac\sigma E,\\\space\\
E=\frac{\sigma L}{L_s-L},\\\space\\
\sigma=T/A,\\\space\\
E=\frac{TL}{A(L_s-L)},\\\space\\
E=199\text{ GPa}."